FRANCK-CONDON CALCULATIONS FOR THE Ã^{1}A_{u}(C_{2h})−X̃^{1}Σ_{g}^{+}(D_{∞h}) TRANSITION OF ACETYLENE

\underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

The Ã − X̃ transition of acetylene was the first electronic transition for which a change of point group was confirmed in detail (1,2). This paper presents harmonic calculations of intensities of the vibrational structure of this transition, using the force field of Tobiason et al. (3) for the Ã state. The Ã − X̃ transition correlates with the forbidden ^{1}Σ_{u}^{−} − ^{1}Σ_{g}^{+} at linear geometries, and the transition moment is assumed to be proportional to the q′′_{4} bending coordinate. There are difficulties in the calculation because the usual harmonic wavefunctions of the upper state do not behave correctly at linear geometries. To correct this, a different calculation would have to be done for each K_{a}-value. Approximate results for the intensity integrals are presented. These agree well with measured values for the v′_{3} − 0 bands in the absorption spectrum with v′_{3} up to 5 (1), but the agreement is only qualitative for recently measured emission bands (4) which peak near v′′_{4} ≈ 14, and must be seriously affected by anharmonicity.

1. C. K. Ingold and G. W. King, J. Chem. Soc., 2702–2755 (1953).

2. K. K. Innes, J. Chem. Phys. 22, 863–876 (1954).

3. J. D. Tobiason, A. L. Utz, E. L. Sibert III, and F. F. Crim, J. Chem. Phys. 99, 5762–5767 (1993).

4. M. P. Jacobson, R. W. Field, et al., personal communication.